OpenStudy (anonymous):

As a fundraiser, a school club is selling posters. The printer charges a $19 set-up fee, plus $0.40 for each poster. Then the cost y in dollars to print x posters is given by the linear equation y=0.40x+19. (a) What is the cost y in dollars to print 50 posters? To print 100 posters? The printing of 50 posters costs $ The printing of 100 posters costs $ (b) Find the number of posters x if the printer billed the club for costs of $91.00. The number of posters, printed for $91.00, is (c) Write the information from parts (a) and (b) as three ordered pairs. The first pair from (a) is The second pair from (a) is The pair from (b) is

OpenStudy (anonymous):

(a) y= 0.40(50) +19~~~~~ (b)91.00= 0.40(x) + 19 ~~~~~